OpenStudy (anonymous):

What is the slope of the line 18x+3y= -162 ? A) -54 B) 6 C)9 D) -6 Need help.

OpenStudy (anonymous):

solve for y. Your slope will be "m" in the formula y = mx+b

OpenStudy (anonymous):

I tried but it comes out incorrect.

OpenStudy (anonymous):

hmm. y = -6x -54 is this what you got?

OpenStudy (anonymous):

Eh no. I had a positive six the first time and then some other numbers.

OpenStudy (anonymous):

well if you take 18x+3y = -162, you first subtract 18x from both sides, and then you'll divide both sides by 3, giving you my answer above. making the slope -6

OpenStudy (anonymous):

18x-3y= -162 -18 -18 equals 3y= -180 and If I divide by 3 I get -60 ._.

OpenStudy (anonymous):

no, it must be 18x there is a variable there you cannot combine 18x and the -162. x is a variable and 18 is multiplied by it, its just like saying 18*x

OpenStudy (anonymous):

Okay. Can you show me step by step on how to find the slope though?

OpenStudy (anonymous):

sure, one second

OpenStudy (anonymous):

\[y = mx+b\] the value 'm' represents slope in this formula. \[18x+3y = -162\] this is your problem, first you subtract 18x from both sides\[3y = -18x-162\] now you divide both sides by 3 to isolate y and put the equation into the formula noted above.\[y = -6x - 54\] as you can see if you say y = mx+b your 'm' value is -6.

OpenStudy (anonymous):

do you understand? if you need clarification just ask.

OpenStudy (anonymous):

I understand now. I didn't do the first step but now I got it. Thanks!

OpenStudy (anonymous):

no problem!